factor this polynomial 2z^2-196c^2
Answer by fractalier(6550)

Well, we start by pulling out a 2...
2z^2 - 196c^2 =
2(z^2 - 98c^2)
Now this is the difference of squares, except that 98 is not a perfect square, but we can still use the proper form, like this
2(z^2 - 98c^2) =
2(z + 7c*sqrt(2))(x - 7c*sqrt(2))